OpenStudy (anonymous):

a salesperson earns a salary of $11,000 a year, she also makes a comission of 20% on her sales... if her sales for the year total was 79,000........how much money was she paid for the year??

OpenStudy (anonymous):

11,000 + 79,000 * 0.2 = 11,000 + 15,800 = 26,800